# appReferences in the Appendix

### v2

@@ Appendix C Main Results
In this section, we demonstrate the proof for all theoretical results that we proposed in the main body of the paper. We instate Assumption 1 and 2 are satisfied.
C.1 Proof of Lemma 3.1
By the boundedness results in Lemma A.1 and Lemma A.2, define
then we complete the proof.
C.2 Proof of Theorem 3.1
@@
    If we use LP-based method on the first batch, then we will have $\Delta_{T_{1}}\leq\log(f)$. Total regret follows
@@
    If we have a warm start for the first batch, with the initialization $\lVert\bm{p}_{0}-\bm{p}^{*}\rVert\leq f^{-1/3}$, then the first batch regret achieves $\Delta_{T_{1}}\in\mathcal{O}(f^{1/3}+\log(f))$ by Lemma B.4. Thus, the total regret follows
@@ Appendix D Auxiliary Results
In this section, we provide some auxiliary results to support the proof in the previous three sections. These lemmas focus on pure mathematical derivations.
@@ D.1 Technical Support for LP-based Analysis
Proof. We analyze each term in (B.3). The key technique we used is to take conditional expectations and simplify the double summations.
@@
    =\ = $\frac{1}{(T-(j+1)f)^{2}}\ \mathbb{E}\left[\sum_{\ell=jf+1}^{(j+1)f}\sum_{s=jf+1}^{(j+1)f}\mathbb{E}\Big{[}(d_{i,jf}-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*}))(d_{i,jf}-{a}_{i,s}I(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{jf}^{*}))\mid\bm{d}_{jf}\Big{]}\right]$
    =\ = $\frac{1}{(T-(j+1)f)^{2}}\ \mathbb{E}\left[\sum_{\ell=jf+1}^{(j+1)f}\mathbb{E}\Big{[}(d_{i,jf}-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*}))^{2}\mid\bm{d}_{jf}\Big{]}\right]$
@@
    =\ = $\frac{f(\bar{a}+\bar{d})^{2}}{(T-(j+1)f)^{2}}\ \ \ (\text{since }\mathbb{E}\Big{[}d_{i,jf}-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})|\bm{d}_{jf}\Big{]}=0\text{ for binding terms})$
@@
    =\ = $\frac{\mathbb{E}\Bigg{[}\sum_{\ell,s=jf+1}^{jk+1)f}\mathbb{E}\Big{[}({a}_{i,\ell}(I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})-I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f})))\cdot({a}_{i,s}(I(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{jf}^{*})-I(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{(j+1)f})))\Big{|}\bm{d}_{jf}\Big{]}\Bigg{]}}{(T-(j+1)f)^{2}}.$
@@
    =\ = $\mathbb{E}\Bigg{[}\mathbb{E}\Big{[}({a}_{i,\ell}(I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})-I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f})))({a}_{i,s}(I(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{jf}^{*})-I(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{(j+1)f})))\mid\bm{a}_{i},\bm{a}_{s}\Big{]}\mid\bm{d}_{jf}\Bigg{]}$
    =\ = $\mathbb{E}\Bigg{[}{a}_{i,\ell}\mathbb{E}\Big{[}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})-I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f})\mid\bm{a}_{\ell}\Big{]}\cdot{a}_{i,s}\mathbb{E}\Big{[}I(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{jf}^{*})-I(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{(j+1)f})\mid\bm{a}_{s}\Big{]}\mid\bm{d}_{jf}\Bigg{]}$
    =\ = $\mathbb{E}\Bigg{[}{a}_{i,\ell}(P(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*}\mid\bm{a}_{\ell})-P(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f}\mid\bm{a}_{\ell}))\cdot{a}_{i,s}(P(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{jf}^{*}\mid\bm{a}_{s})-P(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{(j+1)f}\mid\bm{a}_{s}))\mid\bm{d}_{jf}\Bigg{]}$
@@
[…]
@@
    =\ = $\frac{2}{T-(j+1)f}\ \mathbb{E}\left[\sum_{\ell=jf+1}^{(j+1)f}\mathbb{E}\Big{[}{d}_{i,jf}-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})\mid\bm{d}_{jf}\Big{]}\cdot({d}_{i,jf}-d_{i})\right]$
[…]
@@
    =\ = $\frac{2}{T-(j+1)f}\ \mathbb{E}\Bigg{[}\sum_{\ell=jf+1}^{(j+1)f}({d}_{i,jf}-{d}_{i}){a}_{i,\ell}\ \mathbb{E}\Big{[}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})-I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f})\mid\bm{a}_{\ell}\Big{]}\Bigg{]}$
@@
[…]
    =\ = $\frac{4\mu\bar{a}^{2}\sqrt{C_{lp}}}{(k-j-1)\sqrt{(j+1)f}}\sqrt{\mathbb{E}[({d}_{i,jf}-{d}_{i})^{2}]}.$
@@
[…]
    =\ = $\frac{2\bar{a}(\bar{a}+\bar{d})\sqrt{1+\frac{1}{j}\mu^{2}\bar{a}^{2}C_{lp}}}{(k-j-1)^{2}f}.$
@@
Lemma D.2.
Proof. We consider a general sequence $\{z_{j}\}_{j=1}^{k}$ with
@@ D.2 Technical Support for First-order Analysis
Following the updated rule of first-order method, we derive last dual price satisfies:
@@
Proof. According to Lemma 3.3, take $\alpha_{t}=\alpha<\frac{1}{\nu\lambda}$, we derive the conditional expectation as
@@
By LP-convergence result in Lemma 3.2, we know $\mathbb{E}\lVert\bm{p}_{kf}-\bm{p}^{*}\rVert\leq\frac{C_{lp}}{\sqrt{kf}}$. By Proposition 3.3 in [8], we know the event $E\coloneqq\lVert\bm{p}_{kf}-\bm{p}^{*}\rVert\leq\frac{C_{lp}}{\sqrt{kf}}$ with probability $\mathbb{P}\geq 1-\frac{1}{T^{4}}$. By Lemma 3.1, we know $\lVert\bm{p}_{t}\rVert\leq\bar{p}$.

### v3

@@ Appendix C Main Results
In this section, we demonstrate the proof for all theoretical results that we proposed in the main body of the paper. We instate Assumption 2.1 and 2.2 are satisfied.
C.1 Proof of Theorem 3.1.
@@
    If we use LP-based method on the first batch, then we will have $\Delta_{T_{1}}\leq\log(f)$. Total performance follows
@@
    If we have a warm start for the first batch, with the initialization $\lVert\bm{p}_{0}-\bm{p}^{*}\rVert\leq f^{-1/3}$, then the first batch regret achieves $\Delta_{T_{1}}\in\mathcal{O}(f^{1/3}+\log(f))$ by Lemma B.10. Thus, total performance follows
@@ Appendix D Auxiliary Results
In this section, we provide auxiliary results to support the proof in the previous three sections. These lemmas focus on pure mathematical derivations.
@@ D.1 Technical Support for LP-based Analysis
Lemma D.2.
Proof of ˜D.1.
We analyze each term in (B.3). The key technique we use is to take conditional expectations and simplify the double summations.
@@
    $=\$$\frac{1}{(T-(j+1)f)^{2}}\ \mathbb{E}\left[\sum_{\ell=jf+1}^{(j+1)f}\sum_{s=jf+1}^{(j+1)f}\mathbb{E}\Big[(d_{i,jf}-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*}))(d_{i,jf}-{a}_{i,s}I(r_{s}>\bm{a}_{s}^{\top}\bm{p}_{jf}^{*}))\mid\bm{d}_{jf}\Big]\right]$
    $=\$$\frac{1}{(T-(j+1)f)^{2}}\ \mathbb{E}\left[\sum_{\ell=jf+1}^{(j+1)f}\mathbb{E}\Big[(d_{i,jf}-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*}))^{2}\mid\bm{d}_{jf}\Big]\right]$
@@
    $\leq\$$\frac{(\bar{a}+\bar{d})^{2}}{(k-j-1)^{2}f}.\ \ \ (\text{since }\lfloor T\rfloor=k\cdot f)$
@@
[…]
@@
[…]
@@
    $\leq\$$\frac{1}{(T-(j+1)f)^{2}}\ \mathbb{E}\left[\sum_{\ell=jf+1}^{(j+1)f}\bar{a}^{2}+\sum_{\ell\neq s,\ell=jf+1}^{(j+1)f}\mu^{2}\bar{a}^{4}\ \mathbb{E}\big[(\bm{p}_{(j+1)f}-\bm{p}_{jf}^{*})^{2}\big|\bm{d}_{jf}\big]\Big|\bm{d}_{jf}\right]$
    $\leq\$$\frac{1}{(T-(j+1)f)^{2}}\ \Big(f\bar{a}^{2}+f^{2}\mu^{2}\bar{a}^{4}\mathbb{E}[(\bm{p}_{(j+1)f}-\bm{p}_{jf}^{*})^{2}\big|\bm{d}_{jf}]\Big)$
@@
    $=\$$\frac{2}{T-(j+1)f}\ \mathbb{E}\left[\sum_{\ell=jf+1}^{(j+1)f}\mathbb{E}\Big[{d}_{i,jf}-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})\mid\bm{d}_{jf}\Big]\cdot({d}_{i,jf}-d_{i})\right]$
    $=\$$0.\ \ \ \text{ (by the definition of binding terms)}$
@@
    $=\$$\frac{2}{T-(j+1)f}\ \mathbb{E}\Bigg[\sum_{\ell=jf+1}^{(j+1)f}({d}_{i,jf}-{d}_{i}){a}_{i,\ell}\ \mathbb{E}\Big[I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})-I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f})\mid\bm{a}_{\ell}\Big]\Bigg]$
    $\leq\$$\frac{2}{T-(j+1)f}\sum_{\ell=jf+1}^{(j+1)f}\mathbb{E}\Bigg[({d}_{i,jf}-{d}_{i}){a}_{i,\ell}\ \Big[P(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*}\mid\bm{a}_{\ell})-P(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f}\mid\bm{a}_{\ell})\Big]\Bigg]$
@@
    $\leq\$$\frac{2\mu\bar{a}^{2}}{T-(j+1)f}\sum_{\ell=jf+1}^{(j+1)f}\sqrt{\mathbb{E}[({d}_{i,jf}-{d}_{i})^{2}]}\cdot\sqrt{\mathbb{E}[(\bm{p}_{(j+1)f}-\bm{p}_{jf}^{*})^{2}]}\ \ \ \text{(by Cauchy's inequality)}$
    $\leq\$$\frac{2\mu\bar{a}^{2}}{T-(j+1)f}\cdot\frac{\sqrt{C_{lp}}f}{\sqrt{jf}}\sqrt{\mathbb{E}[({d}_{i,jf}-{d}_{i})^{2}]}$
    $=\$$\frac{4\mu\bar{a}^{2}\sqrt{C_{lp}}}{(k-j-1)\sqrt{(j+1)f}}\sqrt{\mathbb{E}[({d}_{i,jf}-{d}_{i})^{2}]}.$
@@
    $\leq\$$2\sqrt{\mathbb{E}\left(\frac{\sum_{\ell=jf+1}^{(j+1)f}{d}_{i,jf}-{a}_{\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})}{T-(j+1)f}\right)^{2}}\sqrt{\mathbb{E}\left(\frac{\sum_{\ell=jf+1}^{(j+1)f}{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{jf}^{*})-{a}_{i,\ell}I(r_{\ell}>\bm{a}_{\ell}^{\top}\bm{p}_{(j+1)f})}{T-(j+1)f}\right)^{2}}$
@@
    $=\$$\frac{2\bar{a}(\bar{a}+\bar{d})\sqrt{1+\frac{1}{j}\mu^{2}\bar{a}^{2}C_{lp}}}{(k-j-1)^{2}f}.$
@@
Proof of ˜D.2.
We consider a general sequence $\{z_{j}\}_{j=1}^{k}$ with
@@ D.2 Technical Support for First-order Analysis
Following the updated rule of first-order method, we derive the last dual price satisfies:
@@
Proof of ˜D.3.
According to Lemma A.6, take $\alpha_{t}=\alpha<\frac{1}{\nu\lambda}$, we derive the conditional expectation as
@@
By LP-convergence result in Lemma A.4, we know $\mathbb{E}\lVert\bm{p}_{kf}-\bm{p}^{*}\rVert\leq\frac{C_{lp}}{\sqrt{kf}}$. By Proposition 3.3 in (Gao et al., 2024), we know the event $E\coloneqq\lVert\bm{p}_{kf}-\bm{p}^{*}\rVert\leq\frac{C_{lp}}{\sqrt{kf}}$ with probability $\mathbb{P}\geq 1-\frac{1}{T^{4}}$. By Lemma B.2, we know $\lVert\bm{p}_{t}\rVert\leq\bar{p}$.
@@
Appendix E Supplementary Experiments
In this section, we provide more experiments to further evaluate the performance of our algorithms. We consider a more general and complex input distribution and include additional comparisons with recent methods.
E.1 New Distribution
As an extension of Section 4.1, our goal is to evaluate the main algorithms with different choices of re-solving frequency $f\in\{T^{1/3},T^{1/2},T^{2/3}\}$. We consider a more complex distribution for reward and resource consumption requests and guarantee that Assumptions 2.1 and 2.2 are still satisfied.
Consider
Input III: $\ a_{it}\sim\min(1,\max(0,1+z)),\ r_{t}\sim\text{Unif}[0,1]$
$\text{ where }z\sim t(1):\text{Student's t-distribution with 1 degree of freedom}.$
We generate $\{r_{t},\bm{a}_{t}\}_{t=1}^{T}$ from Input III and keep other parameters $T\in[10^{2},10^{6}]$ and $d_{i}\sim\text{Uniform}[1/3,2/3]$ the same in Section 4.1. We report the average result over 100 trials for each experiment and use the classic first-order method with $\mathcal{O}(T^{1/2})$ regret (Algorithm˜4) as a baseline.
Figure 5: Algorithms under New Distribution.
$T$ First-Order Low freq Mid freq High freq
Algorithm 1 $10^{2}$ 3.78 1.68 1.26 1.20
$10^{3}$ 10.69 2.60 2.10 1.54
$10^{4}$ 36.79 4.79 3.67 2.54
$0^{5} 11$7.13 10.45 6.17 3.56
$10^{6}$ 377.84 34.54 14.30 5.48
Algorithm 2 $10^{2}$ 3.78 1.24 1.17 1.13
$10^{3}$ 10.69 2.30 1.96 1.52
$10^{4}$ 36.79 4.12 2.99 2.15
$0^{5} 11$7.13 9.78 4.88 3.05
$10^{6}$ 377.84 25.96 8.90 4.86
Figure 7: Regret for various re-solving frequencies.
Under the new distribution, our algorithms still exhibit a strong regret performance. As shown in Table 7 and Figure 7, we observe that regret decreases as the re-solving frequency increases. This trend holds consistently across both algorithms and is consistent with the guarantees of Theorem 3.2.
E.2 More Comparison
Building on the comparison of baseline methods in Section 4.2, we evaluate our algorithms against recent works. As a reminder, our algorithms employ frequent LP-solving to learn online dual price under continuous support. Li et al. (2024) considers a similar problem using infrequent LP-solving to update the dual variable but under finite support.
Figure 8: Algorithms comparison.
$T$ Regret Algorithm Compute Time (s)
$0^{3} 11$.92 First-Order 0.002
11.12 Infrequent LP-based 0.724
3.77 Algorithm 1 0.790
3.69 Algorithm 2 0.782
$10^{4}$ 36.37 First-Order 0.01
14.20 Infrequent LP-based 0.827
6.69 Algorithm 1 3.015
6.35 Algorithm 2 3.112
$0^{5} 11$0.22 First-Order 0.109
20.90 Infrequent LP-based 1.028
16.41 Algorithm 1 52.95
10.84 Algorithm 2 52.39
$10^{6}$ 312.26 First-Order 1.106
28.80 Infrequent LP-based 4.929
20.58 Algorithm 1 1261.0
14.30 Algorithm 2 1305.4
Figure 9: Regret for various algorithms.
To compare them, we adapt our algorithms to finite support and take the re-solving frequency $f=T^{1/3}$ from Section 4.1; We take the best-performed parameters for the infrequent method where the number of customer types $n=50$ and $\alpha=\beta=0.95$ which control the solving times near start and end. We take the horizon over $T\in[10^{2},10^{6}]$ and report the average result over 100 trials for each experiment. We still use the first-order method in Algorithm˜4 as a baseline.
Table 9 and Figure 9 demonstrate the algorithm regret and computation time across different horizons. While the infrequent LP-based method has faster computation, our algorithms show a competitive performance in decision optimality and achieve lower regret under finite support.